OpenStudy (anonymous):

what is the solution of this linear system 3x+y=-2 y=-x+4

OpenStudy (marihelenh):

Put what y is equal to back into the first equation, so \[3x+(-x+4)=-2\]

OpenStudy (marihelenh):

Combine like terms first. \[3x-x+4=-2\] \[2x+4=-2\] Then you can subtract the 4 from both sides to get x on one side. \[2x=-6\] To finish, just divide both sides by 2 to find what x is equal to.

OpenStudy (marihelenh):

\[x=-3\]

OpenStudy (anonymous):

got it thanks